# S, T, U Parameters in The B-LSSM

### v1

@@ V.1 Neutralino-Chargino sector
$-\sum_{i,j}^{7}(N^{*}_{j3}N_{i3}-N^{*}_{j4}N_{i4})^{2}\}=0$
$div_{T}\sim$ $-\frac{G^{2}}{4m_{Z}^{2}}\{\mathcal{A}_{1}-\mathcal{A}_{2}\}$
@@ V.2 Squark sector
When superquarks are loop particles in the one-loop self-energy diagrams, the cancellation of divergences in the $S$, $T$, and $U$ parameters can be demonstrated as follows.
@@
$\sum^{6}_{k=1}(Z^{U}_{ik}Z^{U*}_{jk})=\sum^{6}_{k=1}(Z^{D}_{ik}Z^{D*}_{jk})=\delta_{ij}$
@@
A study provided the global fitting of $S$ $T$ and $U$ parameters with recently measured top quark mass data from CMS laboratorydeBlas:2022hdk :
$S=0.005\pm 0.096\qquad$ $T=0.04\pm 0.12\qquad U=0.134\pm 0.087$
$\rho_{ST}=0.91\qquad\quad\qquad$ $\rho_{SU}=-0.65\quad\qquad\rho_{TU}=-0.88$
In order to study how the parameters affect the vaules of the $S$, $T$ and $U$ parameters, we use the Mathematica to calculate the oblique parameters of the B-LSSM. Some parmeters we assumed are: $A_{t}=3~{}\rm TeV$, $A_{b}=3~{}\rm TeV$, $g_{YB}=-0.3$, $g_{B}=0.5$, $\tan\beta=20$, $\tan\beta^{\prime}=1.5$, $m_{Z^{\prime}}=4500~{}\rm GeV$, $\mu=200~{}\rm GeV$, $\mu_{B}=800~{}\rm GeV$, $m_{B}=600~{}\rm GeV$, $m_{BL}=800~{}\rm GeV$ $m_{1}=200~{}\rm GeV$ and $m_{2}=300~{}\rm GeV$. We use this set of data as a benchmark and run some of the variables respectively. By filtering, we can obtain some parameters that are sensitive to $S$, $T$ and $U$ parameters
@@
We performed a random three-dimensional sampling of the $S$, $T$, and $U$ parameters in Eq. (94-96) based on the probability distribution results obtained from the experimental fits. Each randomly selected point corresponds to a unique set of values for $\Delta s$, $s^{\prime 2}$, and $g_{B}$. We then conducted statistical analysis on the resulting $\Delta s$, $s^{\prime 2}$, and $g_{B}$ values to generate probability histograms. As shown in Fig. 5, the distribution of $\Delta s$ does not exceed one standard deviation of the current experimental measurements, making it challenging to extract further useful information from the measurement of the Weinberg angle. Additionally, the probability distribution of $g_{B}$ exhibits a sharp peak around $g_{B}=0.09$, with a rapid decay for $g_{B}>0.2$. Analysis reveals that the probability for $0.210.03<g_{B}<$ 0.21 exceeds ninety percent when compared to the entire range.
@@ VII Conclusions
Using the effective Lagrangian and experimental measurements, the coupling constant of the $U(1)$ group in models with an additional $U(1)$ symmetry can be related to the $S$, $T$, and $U$ parameters. Through analysis, effective constraints on these coefficients can be derived.
By comparing the effective Lagrangian with the model’s Lagrangian and combining the results from fitting experimental measurements, the coupling constant of the $U(1)$ group in models with an additional $U(1)$ symmetry can be related to the $S$, $T$, and $U$ parameters. Through analysis, effective constraints on these coefficients can be imposed.
@@ Acknowledgements.
The work was supported by the Natural Science Foundation of Guangxi Autonomous Region with Grant No. 2022GXNSFDA035068.
@@ Appendix A Integral formulas for some Feynman diagrams
Here we provide some integral formulas for vacuum polarization diagrams, we will unify them here in the form of $A$ or $B$ functions. We define $A$ fuctions:
$i\pi^{2}A_{0}(m)=$ $\mu^{4-n}\int d^{n}q\frac{1}{q^{2}+m^{2}-i\epsilon}$
$i\pi^{2}A_{\mu}(m)=$ $\mu^{4-n}\int d^{n}q\frac{q_{\mu}}{q^{2}+m^{2}-i\epsilon}=0$
$i\pi^{2}A_{\mu\nu}(m)=$ $\mu^{4-n}\int d^{n}q\frac{q_{\mu}q_{\nu}}{q^{2}+m^{2}-i\epsilon}=i\pi^{2}A_{2}(m)g_{\mu\nu}$ (101)
@@
$i\pi^{2}B_{0}(p^{2};m_{1},m_{2})=$ $\mu^{4-n}\int d^{n}q\frac{1}{d_{0}d_{1}}$
$i\pi^{2}B_{\mu}(p^{2};m_{1},m_{2})=$ $\mu^{4-n}\int d^{n}q\frac{q_{\mu}}{d_{0}d_{1}}=i\pi^{2}B_{1}(p^{2};m_{1},m_{2})p_{\mu}$
$i\pi^{2}B_{\mu\nu}(p^{2};m_{1},m_{2})=$ $\mu^{4-n}\int d^{n}q\frac{q_{\mu}q_{\nu}}{d_{0}d_{1}}=i\pi^{2}[B_{21}(p^{2};m_{1},m_{2})p_{\mu}p_{\nu}+B_{22}(p^{2};m_{1},m_{2})g_{\mu\nu}]$ (102)
@@
$A_{0}(m)\sim-m^{2}$ $A_{2}(m)\sim\frac{m^{4}}{4}$ $B_{0}(p^{2};m_{1},m_{2})\sim 1$
$B_{1}(p^{2};m_{1},m_{2})\sim-\frac{1}{2}$ $B_{21}(p^{2};m_{1},m_{2})\sim\frac{1}{3}$ $B_{22}(p^{2};m_{1},m_{2})\sim-\frac{m^{2}_{1}+m_{2}^{2}}{4}-\frac{p^{2}}{12}$
@@
$\mathcal{M}_{a}=$ $-\mathcal{C}_{a}\{4B_{22}(p^{2};m_{1},m_{2})g_{\mu\nu}+[B_{0}+4B_{1}+4B_{21}](p^{2};m_{1},m_{2})p_{\mu}p_{\nu}\}$
$\mathcal{M}_{b}=$ $\mathcal{C}_{b}A_{0}(m)g_{\mu\nu}$ (103)
$\mathcal{C}_{a}$ and $\mathcal{C}_{b}$ are the coefficients from the vertexs. We expand and sum the coefficients of the involved graphs.
$\mathcal{C}_{a}^{WW}=$ $-\frac{1}{2}g^{2}_{2}\sum^{3}_{k=1}(Z^{D*}_{ik}Z^{U}_{jk})\sum^{3}_{l=1}(Z^{U*}_{jl}Z^{D}_{il})$
$\mathcal{C}_{b}^{WWD}=$ $\frac{1}{2}g^{2}_{2}\sum^{3}_{k=1}(Z^{D*}_{ik}Z^{D}_{ik})\qquad\mathcal{C}_{b}^{WWU}=\frac{1}{2}g^{2}_{2}\sum^{3}_{k=1}(Z^{U*}_{jk}Z^{U}_{jk})$
$\mathcal{C}_{a}^{\gamma\gamma U}=$ $-\frac{4}{9}e^{2}(\sum_{k=1}^{3}Z^{U*}_{ik}Z^{U}_{jk}+\sum_{k=4}^{6}Z^{U*}_{ik}Z^{U}_{jk})^{2}$
$\mathcal{C}_{a}^{\gamma\gamma D}=$ $-\frac{1}{9}e^{2}(\sum_{k=1}^{3}Z^{D*}_{ik}Z^{D}_{jk}+\sum_{k=4}^{6}Z^{D*}_{ik}Z^{D}_{jk})^{2}$
$\mathcal{C}_{b}^{\gamma\gamma U}=$ $\frac{8}{9}e^{2}(\sum_{k=1}^{3}Z^{U*}_{ik}Z^{U}_{ik}+\sum_{k=4}^{6}Z^{U*}_{ik}Z^{U}_{ik})$
$\mathcal{C}_{b}^{\gamma\gamma D}=$ $\frac{1}{18}e^{2}(\sum_{k=1}^{3}Z^{D*}_{jk}Z^{D}_{jk}+\sum_{k=4}^{6}Z^{D*}_{jk}Z^{D}_{jk})$
$\widetilde{\mathcal{C}}_{a}^{ZZU}=$ $-\frac{1}{36}[G(1+2c_{2W})\sum^{3}_{k=1}Z^{U*}_{ik}Z^{U}_{jk}-2G(1-c_{2W})\sum^{6}_{k=4}Z^{U*}_{ik}Z^{U}_{jk}]^{2}$
$\widetilde{\mathcal{C}}_{a}^{ZZD}=$ $-\frac{1}{36}[G(-2-c_{2W})\sum^{3}_{k=1}Z^{D*}_{ik}Z^{D}_{jk}+G(1-c_{2W})\sum^{6}_{k=4}Z^{D*}_{ik}Z^{D}_{jk}]^{2}$
$\widetilde{\mathcal{C}}_{b}^{ZZU}=$ $\frac{1}{18}\{[G(1+2c_{2W})\sum^{3}_{k=1}Z^{U*}_{ik}Z^{U}_{ik}]^{2}+[2G(1-c_{2W})\sum^{6}_{k=4}Z^{U*}_{ik}Z^{U}_{ik}]^{2}\}$
$\widetilde{\mathcal{C}}_{b}^{ZZD}=$ $\frac{1}{18}\{[G(-2-c_{2W})\sum^{3}_{k=1}Z^{D*}_{jk}Z^{D}_{jk}]^{2}+[G(1-c_{2W})\sum^{6}_{k=4}Z^{D*}_{jk}Z^{D}_{jk}]^{2}\}$
$\widetilde{\mathcal{C}}_{a}^{Z\gamma U}=$ $-\frac{e}{9}(\sum_{k=1}^{3}Z^{U*}_{ik}Z^{U}_{jk}+\sum_{k=4}^{6}Z^{U*}_{ik}Z^{U}_{jk})[G(1+2c_{2W})\sum^{3}_{k=1}Z^{U*}_{ik}Z^{U}_{jk}-2G(1-c_{2W})\sum^{6}_{k=4}Z^{U*}_{ik}Z^{U}_{jk}]^{2}$
$\widetilde{\mathcal{C}}_{a}^{Z\gamma D}=$ $-\frac{e}{18}(\sum_{k=1}^{3}Z^{D*}_{ik}Z^{D}_{jk}+\sum_{k=4}^{6}Z^{D*}_{ik}Z^{D}_{jk})[G(-2-c_{2W})\sum^{3}_{k=1}Z^{D*}_{ik}Z^{D}_{jk}+G(1-c_{2W})\sum^{6}_{k=4}Z^{D*}_{ik}Z^{D}_{jk}]^{2}$
$\widetilde{\mathcal{C}}_{b}^{Z\gamma U}=$ $\frac{2}{9}e[G(1+2c_{2W})(\sum^{3}_{k=1}Z^{U*}_{ik}Z^{U}_{ik})^{2}-2G(1-c_{2W})(\sum^{6}_{k=4}Z^{U*}_{ik}Z^{U}_{ik})^{2}]$
$\widetilde{\mathcal{C}}_{b}^{Z\gamma D}=$ $\frac{1}{9}e[G(-2-c_{2W})(\sum^{3}_{k=1}Z^{D*}_{jk}Z^{D}_{jk})^{2}+G(1-c_{2W})(\sum^{6}_{k=4}Z^{D*}_{jk}Z^{D}_{jk})^{2}]$
@@
$\widetilde{\mathcal{C}}^{ZZ}=$ ${c^{\prime}}^{2}\mathcal{C}^{ZZ}-2c^{\prime}s^{\prime}\mathcal{C}^{ZZ^{\prime}}+{s^{\prime}}^{2}\mathcal{C}^{Z^{\prime}Z^{\prime}}$
$\widetilde{\mathcal{C}}^{Z\gamma}=$ $c^{\prime}\mathcal{C}^{Z\gamma}-s^{\prime}\mathcal{C}^{Z^{\prime}\gamma}$ (104)
Appendix B Slef-Energys of Charginos and Neutralinos

### v2

@@ V.1 Neutralino-Chargino sector
$-\sum_{i,j}^{7}(N^{*}_{j3}N_{i3}-N^{*}_{j4}N_{i4})^{2}\}=0.$
$div_{T}\sim$ $-\frac{G^{2}}{4m_{Z}^{2}}\{\mathcal{A}_{1}-\mathcal{A}_{2}\}.$
@@ V.2 Squark sector
When squarks are loop particles in the one-loop self-energy diagrams, the cancellation of divergences in the $S$, $T$, and $U$ parameters can be demonstrated as follows.
@@
$\sum^{6}_{k=1}(Z^{U}_{ik}Z^{U*}_{jk})=\sum^{6}_{k=1}(Z^{D}_{ik}Z^{D*}_{jk})=\delta_{ij}.$
This work focuses on the neutralino-chargino and squark sectors of the B-LSSM, which are expected to provide the dominant new-physics contributions to electroweak precision observables, and within this framework we have constructed and verified the gauge-invariant formalism. Although the Higgs sector involves non-Abelian symmetry breaking and its mixing with Goldstone bosons, which introduces specific complexities in maintaining gauge invariance, its loop contributions to the $S$, $T$, and $U$ parameters have been shown to be extremely small already in the Standard Model. In the B-LSSM, the Higgs fields associated with the additional $U(1)$ breaking typically acquire masses well above the electroweak scale, further suppressing their contributions. Moreover, for the Higgs sector to yield a significant effect, it would generally require mixing between different generations of scalar fields, and such mixing is usually suppressed by the large mass scales. Therefore, at this stage of systematically studying the main sources of new-physics contributions, we have not included the Higgs sector in the detailed calculations.
@@
A global fit of the $S$, $T$, and $U$ parameters incorporating recent top quark mass measurements from CMS has been provided in Ref.ParticleDataGroup:2024cfk :
$S=-0.04\pm 0.10,\qquad$ $T=0.01\pm 0.12,\qquad U=-0.01\pm 0.09,$
$\rho_{ST}=0.93,\qquad\quad\qquad$ $\rho_{SU}=-0.87,\quad\qquad\rho_{TU}=-0.70.$
In order to study how the parameters affect the vaules of the $S$, $T$ and $U$ parameters, we use the Mathematica to calculate the oblique parameters of the B-LSSM. Some parmeters we assumed are: $A_{t}=3~\rm TeV$, $A_{b}=3~\rm TeV$, $g_{YB}=-0.3$, $g_{B}=0.5$, $\tan\beta=20$, $\tan\beta^{\prime}=1.5$, $m_{Z^{\prime}}=4500~\rm GeV$, $\mu=200~\rm GeV$, $\mu_{B}=800~\rm GeV$, $m_{B}=600~\rm GeV$, $m_{BL}=800~\rm GeV$ $m_{1}=200~\rm GeV$ and $m_{2}=300~\rm GeV$. We use this set of data as a benchmark and run some of the variables respectively. Through a parameter scan, we can identify which model parameters are most sensitive to the $S$, $T$, and $U$ observables.
@@
Figure 5: The dependence of the oblique parameters $S,T$, and $U$ on the B-LSSM distinctive parameters $g_{YB}$, $g_{B}$, and $m_{Z^{\prime}}$. The SUSY-breaking parameters are fixed at the benchmark point.
To genuinely highlight the distinctive features of the B-LSSM as opposed to the MSSM, it is essential to illustrate how the extended gauge sector parameters influence the electroweak precision observables. In Fig. 5, we present the dependence of the $S$, $T$, and $U$ parameters on the unique B-LSSM parameters: the gauge kinetic mixing coupling $g_{YB}$, the $U(1)_{B-L}$ gauge coupling $g_{B}$, and the extra gauge boson mass $m_{Z^{\prime}}$, while keeping other supersymmetric parameters fixed at the aforementioned benchmark point.
As shown in the panels of Fig. 5, varying $m_{Z^{\prime}}$ from 3000 GeV to 6000 GeV leads to a smooth, monotonic modulation of the oblique parameters. Because the $Z^{\prime}$ boson is heavy, its decoupling behavior ensures that the absolute variations remain strictly perturbative. Furthermore, the figures demonstrate the pronounced effects of the extra gauge couplings. While $g_{B}$ dictates the interaction strength of the $U(1)_{B-L}$ sector, the parameter $g_{YB}$ explicitly governs the gauge kinetic mixing between the $U(1)_{Y}$ and $U(1)_{B-L}$ groups. This kinetic mixing effect directly modifies the $Z-Z^{\prime}$ mass matrix, systematically shifting the $T$ parameter closer to zero while slightly reducing the $S$ parameter. These results explicitly demonstrate that while the loop-level sparticle mass patterns set the baseline, the unique $U(1)_{B-L}$ extensions provide a finite and distinctive tree-level modulation.
We performed a random three-dimensional sampling of the $S$, $T$, and $U$ parameters in Eq. (43-45) based on the probability distribution results obtained from the experimental fits. Each randomly selected point corresponds to a unique set of values for $\Delta s$, $s^{\prime 2}$, and $g_{B}$. We then conducted statistical analysis on the resulting $\Delta s$, $s^{\prime 2}$, and $g_{B}$ values to generate probability histograms. As shown in Fig. 6, the distribution of $\Delta s$ does not exceed one standard deviation of the current experimental measurements, making it challenging to extract further useful information from the measurement of the Weinberg angle. Additionally, the probability distribution of $g_{B}$ exhibits a sharp peak around $g_{B}=0.08$, with a rapid decay for $g_{B}>0.05$. Analysis reveals that the probability for $0.05<g_{B}<0.2$ exceeds ninety percent when compared to the entire range.
@@ VII Conclusions
By comparing the effective Lagrangian with the model Lagrangian and incorporating constraints from experimental fits, the coupling constant of the additional $U(1)$ group can be related to the $S$, $T$, and $U$ parameters. This analysis allows us to derive effective constraints on these coefficients.
Finally, comparing the loop-level supersymmetric contributions with the tree-level extra gauge effects reveals a distinct physical hierarchy within the B-LSSM. Under current experimental precision, the globally fitted electroweak observables impose significantly more stringent constraints on the additional $U(1)_{B-L}$ gauge parameters (e.g., $g_{B}$ and $m_{Z^{\prime}}$) than on the detailed supersymmetric mass spectrum. The latter, entering primarily via loop corrections, is naturally suppressed by the decoupling of heavy sparticles. In contrast, the extra gauge sector directly modulates the observables through tree-level kinetic mixing and mass matrix deformations. Consequently, contemporary precision measurements act as a powerful and direct probe of the extended gauge sector, validating the necessity of investigating the distinctive gauge extensions beyond the minimal supersymmetric framework.
@@ Acknowledgements.
The work was supported by the Natural Science Foundation of Guangxi Autonomous Region with Grant No. 2022GXNSFDA035068.The Natural Science Foundation of Hebei Province under Grant No.A2022104001,the Foundation of Baoding University under Grant No. 2023Z01
@@ Appendix A Integral formulas for some Feynman diagrams
We provide the integral formulas for the vacuum polarization diagrams, expressing them uniformly in terms of the standard Passarino-Veltman A and B functions. We define $A$ fuctions:
$i\pi^{2}A_{0}(m)=$ $\mu^{4-n}\int d^{n}q\frac{1}{q^{2}+m^{2}-i\epsilon},$
$i\pi^{2}A_{\mu}(m)=$ $\mu^{4-n}\int d^{n}q\frac{q_{\mu}}{q^{2}+m^{2}-i\epsilon}=0,$
$i\pi^{2}A_{\mu\nu}(m)=$ $\mu^{4-n}\int d^{n}q\frac{q_{\mu}q_{\nu}}{q^{2}+m^{2}-i\epsilon}=i\pi^{2}A_{2}(m)g_{\mu\nu},$ (50)
@@
$i\pi^{2}B_{0}(p^{2};m_{1},m_{2})=$ $\mu^{4-n}\int d^{n}q\frac{1}{d_{0}d_{1}},$
$i\pi^{2}B_{\mu}(p^{2};m_{1},m_{2})=$ $\mu^{4-n}\int d^{n}q\frac{q_{\mu}}{d_{0}d_{1}}=i\pi^{2}B_{1}(p^{2};m_{1},m_{2})p_{\mu},$
$i\pi^{2}B_{\mu\nu}(p^{2};m_{1},m_{2})=$ $\mu^{4-n}\int d^{n}q\frac{q_{\mu}q_{\nu}}{d_{0}d_{1}}=i\pi^{2}[B_{21}(p^{2};m_{1},m_{2})p_{\mu}p_{\nu}+B_{22}(p^{2};m_{1},m_{2})g_{\mu\nu}].$ (51)
@@
$A_{0}(m)\sim-m^{2},$ $A_{2}(m)\sim\frac{m^{4}}{4},$ $B_{0}(p^{2};m_{1},m_{2})\sim 1,$
$B_{1}(p^{2};m_{1},m_{2})\sim-\frac{1}{2},$ $B_{21}(p^{2};m_{1},m_{2})\sim\frac{1}{3},$ $B_{22}(p^{2};m_{1},m_{2})\sim-\frac{m^{2}_{1}+m_{2}^{2}}{4}-\frac{p^{2}}{12}.$
@@
$\mathcal{M}_{a}=$ $-\mathcal{C}_{a}\{4B_{22}(p^{2};m_{1},m_{2})g_{\mu\nu}+[B_{0}+4B_{1}+4B_{21}](p^{2};m_{1},m_{2})p_{\mu}p_{\nu}\},$
$\mathcal{M}_{b}=$ $\mathcal{C}_{b}A_{0}(m)g_{\mu\nu}.$ (52)
$\mathcal{C}_{a}$ and $\mathcal{C}_{b}$ are the coefficients from the vertices. We expand and sum the coefficients of the involved graphs.
$\mathcal{C}_{a}^{WW}=$ $-\frac{1}{2}g^{2}_{2}\sum^{3}_{k=1}(Z^{D*}_{ik}Z^{U}_{jk})\sum^{3}_{l=1}(Z^{U*}_{jl}Z^{D}_{il}),$
$\mathcal{C}_{b}^{WWD}=$ $\frac{1}{2}g^{2}_{2}\sum^{3}_{k=1}(Z^{D*}_{ik}Z^{D}_{ik})\qquad\mathcal{C}_{b}^{WWU}=\frac{1}{2}g^{2}_{2}\sum^{3}_{k=1}(Z^{U*}_{jk}Z^{U}_{jk}),$
$\mathcal{C}_{a}^{\gamma\gamma U}=$ $-\frac{4}{9}e^{2}(\sum_{k=1}^{3}Z^{U*}_{ik}Z^{U}_{jk}+\sum_{k=4}^{6}Z^{U*}_{ik}Z^{U}_{jk})^{2},$
$\mathcal{C}_{a}^{\gamma\gamma D}=$ $-\frac{1}{9}e^{2}(\sum_{k=1}^{3}Z^{D*}_{ik}Z^{D}_{jk}+\sum_{k=4}^{6}Z^{D*}_{ik}Z^{D}_{jk})^{2},$
$\mathcal{C}_{b}^{\gamma\gamma U}=$ $\frac{8}{9}e^{2}(\sum_{k=1}^{3}Z^{U*}_{ik}Z^{U}_{ik}+\sum_{k=4}^{6}Z^{U*}_{ik}Z^{U}_{ik}),$
$\mathcal{C}_{b}^{\gamma\gamma D}=$ $\frac{1}{18}e^{2}(\sum_{k=1}^{3}Z^{D*}_{jk}Z^{D}_{jk}+\sum_{k=4}^{6}Z^{D*}_{jk}Z^{D}_{jk}),$
$\widetilde{\mathcal{C}}_{a}^{ZZU}=$ $-\frac{1}{36}[G(1+2c_{2W})\sum^{3}_{k=1}Z^{U*}_{ik}Z^{U}_{jk}-2G(1-c_{2W})\sum^{6}_{k=4}Z^{U*}_{ik}Z^{U}_{jk}]^{2},$
$\widetilde{\mathcal{C}}_{a}^{ZZD}=$ $-\frac{1}{36}[G(-2-c_{2W})\sum^{3}_{k=1}Z^{D*}_{ik}Z^{D}_{jk}+G(1-c_{2W})\sum^{6}_{k=4}Z^{D*}_{ik}Z^{D}_{jk}]^{2},$
$\widetilde{\mathcal{C}}_{b}^{ZZU}=$ $\frac{1}{18}\{[G(1+2c_{2W})\sum^{3}_{k=1}Z^{U*}_{ik}Z^{U}_{ik}]^{2}+[2G(1-c_{2W})\sum^{6}_{k=4}Z^{U*}_{ik}Z^{U}_{ik}]^{2}\},$
$\widetilde{\mathcal{C}}_{b}^{ZZD}=$ $\frac{1}{18}\{[G(-2-c_{2W})\sum^{3}_{k=1}Z^{D*}_{jk}Z^{D}_{jk}]^{2}+[G(1-c_{2W})\sum^{6}_{k=4}Z^{D*}_{jk}Z^{D}_{jk}]^{2}\},$
$\widetilde{\mathcal{C}}_{a}^{Z\gamma U}=$ $-\frac{e}{9}(\sum_{k=1}^{3}Z^{U*}_{ik}Z^{U}_{jk}+\sum_{k=4}^{6}Z^{U*}_{ik}Z^{U}_{jk})[G(1+2c_{2W})\sum^{3}_{k=1}Z^{U*}_{ik}Z^{U}_{jk}-2G(1-c_{2W})\sum^{6}_{k=4}Z^{U*}_{ik}Z^{U}_{jk}]^{2},$
$\widetilde{\mathcal{C}}_{a}^{Z\gamma D}=$ $-\frac{e}{18}(\sum_{k=1}^{3}Z^{D*}_{ik}Z^{D}_{jk}+\sum_{k=4}^{6}Z^{D*}_{ik}Z^{D}_{jk})[G(-2-c_{2W})\sum^{3}_{k=1}Z^{D*}_{ik}Z^{D}_{jk}+G(1-c_{2W})\sum^{6}_{k=4}Z^{D*}_{ik}Z^{D}_{jk}]^{2},$
$\widetilde{\mathcal{C}}_{b}^{Z\gamma U}=$ $\frac{2}{9}e[G(1+2c_{2W})(\sum^{3}_{k=1}Z^{U*}_{ik}Z^{U}_{ik})^{2}-2G(1-c_{2W})(\sum^{6}_{k=4}Z^{U*}_{ik}Z^{U}_{ik})^{2}],$
$\widetilde{\mathcal{C}}_{b}^{Z\gamma D}=$ $\frac{1}{9}e[G(-2-c_{2W})(\sum^{3}_{k=1}Z^{D*}_{jk}Z^{D}_{jk})^{2}+G(1-c_{2W})(\sum^{6}_{k=4}Z^{D*}_{jk}Z^{D}_{jk})^{2}].$
@@
$\widetilde{\mathcal{C}}^{ZZ}=$ ${c^{\prime}}^{2}\mathcal{C}^{ZZ}-2c^{\prime}s^{\prime}\mathcal{C}^{ZZ^{\prime}}+{s^{\prime}}^{2}\mathcal{C}^{Z^{\prime}Z^{\prime}},$
$\widetilde{\mathcal{C}}^{Z\gamma}=$ $c^{\prime}\mathcal{C}^{Z\gamma}-s^{\prime}\mathcal{C}^{Z^{\prime}\gamma},$ (53)
Appendix B SELF-ENERGIES of Charginos and Neutralinos